Circular No. 7114

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999ac IN NGC 6063 M. Modjaz, J. Y. King, M. Papenkova, A. Friedman, R. A. Johnson,

W. D. Li, R. R. Treffers, and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova in NGC 6063 during the course of the Lick Observatory Supernova Search with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). SN 1999ac was discovered and confirmed on unfiltered observations taken on Feb. 26.5 (mag ∼ 15.2) and 27.5 UT (mag ∼ 15.0). The new object is located at α = 16^{h}07^{m}15^{s}.01, δ = +7^{o}58′20′′.4 (equinox 2000.0), which is ∼ 23′′.9 east and 29′′.8 south of the nucleus of NGC 6063. KAIT images of the same field on 1999 Feb. 20.5 (limiting mag ∼ 17.5) and 1998 July 21.2 (limiting mag ∼ 19.0) show nothing at the position of the supernova.

XTE J1550–564 R. Jain and C. Bailyn, Yale University; J. McClintock and G. Sobczak,

Harvard-Smithsonian Center for Astrophysics; R. Remillard, Massachusetts Institute of Technology; and J. Orosz, Pennsylvania State University, report YALO consortium observations of XTE J1550–564: “We have obtained B-, V -, and I-band images virtually every night since Jan. 4.3 UT and most recently covering 3.5 hr of elapsed time per night. We find that, between Jan. 4.3 and 22.3, the source was steady at V = 16.04, B = 17.48, and I = 14.15. Then on Jan. 22.3, there was a 0.3-mag dip in the B, V, and I magnitudes. The source then remained steady, with a slight decline to a new level (V = 16.17, B = 17.62, I = 14.29) until Feb. 6.3, when another small dip (0.2 mag in I) occurred. This was followed by another period of steady slow decline. Most recently, a much larger decrease in optical brightness has occurred: since Feb. 21, the source has declined by > 0.5 magnitudes in B, V, and I. RXTE PCA observations show that the 2– 20-keV x-ray flux increased from 6.5 × 10^{−8} erg cm^{−2} s^{−1} on Feb. 21.5 to 7.4 × 10^{−8} erg cm^{−2} s^{−1} on Feb. 23.9. The x-ray spectrum from 2 to 20 keV is dominated by the disk component, but the change in optical intensity has been accompanied by a factor of two increase in the powerlaw component of the x-ray flux from Feb. 21.5 to 23.9. We will continue to monitor this object on a daily basis for the foreseeable future, and we strongly encourage other observations, particularly of optical spectroscopy, to follow the current change in state of this source.”

1999 February 27  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green